OpenStudy (anonymous):

4P(s) + 5O2(g)-> P4O10(s) If .489 mol of P burns what mass of oxygen is used? What mass of P4O10 is produced? answer is 19.6...i just need to know how to write the steps correctly..

OpenStudy (anonymous):

you have equation so start by writing what you have and what you need... n(O2)/n(P) = 5/4 so now you see: n(O2) = 5/4 * n(P) got it?